OpenStudy (anonymous):

1.Write the equation of the line which passes through (5, –2) and is parallel to x = 4.

OpenStudy (anonymous):

\[x=4\] is vertical line. So, the line must be vertical and pass through (5, -2). So, the answer is x=5.

OpenStudy (waleed_imtiaz):

As the line is parallel to y-axis...... means perpendicular to x-axis ..... so its slope will be tan90=infinity..... apply point slope form..... y-y1=m(x-x1) 0=1(x-5); x=+5